OpenStudy (anonymous):

slope of theses points (7,-6) and (7,-9)

OpenStudy (nurali):

m=y2-y1/x2-x1 m=-9+6/7-7 m=-3/0 m=infinte

OpenStudy (anonymous):

does that mean its undifined?

OpenStudy (nurali):

yes

OpenStudy (anonymous):

ok thank you!